# Towards Pervasive Distributed Agentic Generative AI - A State of The Art

### v1

@@ 2.2.3. Planning
The planning module is essential for autonomous agents, enabling them to formulate future action sequences to achieve defined goals (Wang et al., 2024c). It allows agents to move in goal-directed, multi-step problem-solving. The planning module combines environmental feedback, current state, desired outcomes, memory, and profile information to construct a trajectory of actions that effects a transition from the agent’s current state to a target goal state. Various methodologies have emerged for incorporating LLMs into the planning process, each representing a different approach to leveraging their capabilities:
@@
  • LLM-as-Planner directly employs the inherent reasoning abilities of LLMs to generate plans from natural language instructions (Li et al., 2024a). This paradigm emphasizes the autonomous planning capacity of LLMs, relying on their ability to interpret and translate natural language directives into actionable plans.
@@
  • LLM-as-Facilitator uses LLMs to augment existing planning algorithms, such as classical symbolic planners. In this context, LLMs may serve as translators, converting natural language problem descriptions into formal planning languages, such as the Planning Domain Definition Language (PDDL), which are then processed by external planners (Li et al., 2024c; Wang et al., 2024c).
@@
  • Multi-Agent Planning involves the coordination of plans and actions among multiple agents to achieve a shared objective (Han et al., 2024). This necessitates the implementation of mechanisms for inter-agent communication, negotiation, and belief revision. Collaborative planning approaches are exemplified by frameworks such as PlanGEN (Parmar et al., 2025), which employs specialised LLM agents for constraint checking, verification, and selection, and Master (Gan et al., 2025), which proposes a hierarchical multi-agent framework that dynamically generates collaborating agents, validates reasoning, and adjusts confidence-based scoring using Monte Carlo Tree Search (MCTS) to enhance accuracy and efficiency.
@@
One key aspect of the planning module is task decomposition, where complex tasks are broken down into smaller, more tractable sub-problems, facilitating efficient execution. For example, PlanGEN (Parmar et al., 2025) is specifically engineered to augment LLMs’ capacity to generate effective natural language plans through this decomposition process. Similarly, WebPilot’s Planner module (Zhang et al., 2024e) initiates its operation by partitioning complex web tasks into manageable subtasks, thereby constructing a flexible, high-level plan that can adapt to the inherent uncertainties of web environments. Consequently, determining the optimal sequence of these subtasks or individual actions, becomes crucial for achieving the overarching goal. In more complex scenarios, the planning module may also necessitate the allocation of available resources across various planning components (Li et al., 2024a). This is exemplified by the TravelPlanner benchmark (Xie et al., 2024), which evaluates agents’ ability to generate travel itineraries from user queries, demanding the navigation of extensive online resources using specialised tools and adherence to user constraints. This benchmark assesses the agent’s resource allocation across diverse tasks, including city searches, flight bookings, accommodation selection, and dining arrangements. In the current literature planning module employs diverse strategies in task decomposition:
@@
  • Single-Path Planning: The agent follows one trajectory of thought and action at a time, without exploring alternative possibilities. Examples include Chain of Thought prompting, where the LLM reasons step-by-step along a linear path (Wang et al., 2024c).
@@
  • Tree-Based Planning: The agent explores multiple potential thought trajectories, organizing them into a tree structure. This allows for backtracking and evaluating different options before committing to a plan. Techniques like Tree of Thoughts and methods using Monte Carlo Tree Search (MCTS) fall under this category (Li et al., 2024a).
@@
  • Hierarchical Planning: Involves planning at different levels of abstraction, with high-level plans broken down into more detailed, low-level actions or sub-goals (Xi et al., 2025; Zhang et al., 2024e).
@@
Effective integration of feedback within the planning module significantly enhances the ability of LLM-based agents to operate in complex and dynamic environments (Han et al., 2024). This capability allows agents to move beyond static, pre-defined plans and engage in a more adaptive and iterative problem-solving process. So we can categorise planning approaches in 2 groups:
@@
  • Planning without Feedback: The agent formulates a plan upfront and executes it sequentially without adjusting based on intermediate outcomes (Han et al., 2024). This approach is appropriate for tasks of lower complexity, including conversational interaction.
@@
  • Planning with Feedback: The agent receives information from the environment, humans, or other models about the progress and outcomes of its actions (Li et al., 2024a). This feedback is then used to dynamically adjust and refine the plan. Environmental Feedback can include task completion signals or observations after taking an action (Yao et al., 2023). Human feedback can provide guidance or corrections (Franciscatto et al., 2025). Model-based feedback, such as self-reflection or critique from another AI model, can also be valuable (Shinn et al., 2023).
@@
Finally, careful consideration of memory and profiling is vital for effective planning. The memory module provides the historical context, past experiences, and relevant knowledge needed to formulate informed plans. For instance, past successes or failures in similar situations, user preferences, or environmental observations stored in memory can guide the planning process (Shinn et al., 2023). The agent’s profile, defined by the profiling module, can also influence the style and focus of the planning process. For example, an agent profiled as risk-averse might prioritise plans with higher certainty of success, while a creative problem-solver profile might encourage the exploration of novel or less conventional plans (Zhang et al., 2024d).
@@ 2.2.4. Action
The primary goal of the Action Module is to translate the agent’s formulated plans into specific outcomes within the environment. It acts as the execution engine, directly engaging with the surrounding world (Wang et al., 2024c). The intended outcomes of agent actions are diverse and intrinsically linked to the agent’s designated task. Actions may be directed towards achieving concrete objectives, such as item crafting in a game (Wang et al., 2023) or software development task completion (Wang et al., 2024b), where each action contributes directly to the final goal. In the context of a web browsing agent (Lai et al., 2024), actions can include navigating or interacting to specific pages. Actions may also facilitate information sharing and collaboration with other agents (Gan et al., 2025) or human users  (Franciscatto et al., 2025), including conversational exchanges and feedback provision. Furthermore, agents perform exploratory actions to gain new knowledge and expand perception (Wen et al., 2024a), optimizing learning and performance through exploration-exploitation. The Action Module receives from the Planning Module a sequence of steps or a specific action to be taken and translates them into executable actions (Wang et al., 2024c). The Action Module’s operational mechanism varies with the agent’s architecture and action space, which collectively define the agent’s interaction capabilities. Actions can be broadly categorised as follows:
@@
  • External Tool and API Interaction: LLM-driven agents frequently generate natural language commands or structured calls to external tools and APIs (Schick et al., 2023). This allows them to interact with the external world, using resources like search engines for web navigation (Lai et al., 2024) (including clicking, form filling, and scrolling), databases via SQL queries (Shen et al., 2024), code execution environments (Wang et al., 2024b), PDDL solver execution (Chang et al., 2024), or specific application such as FlightSearch API with correctly formatted parameters (Xie et al., 2024). To simplify the integration of diverse APIs, the Model Context Protocol (MCP) (Anthropic, 2024) standardises API access, eliminating manual tracking and description. MCP’s Host-Client-Server architecture facilitates dynamic tool discovery and execution, enabling any MCP-compatible LLM application to use connected server tools.
@@
  • Internal Response Generation: This category includes actions that generate internal responses, such as providing explanations or refining plans (Yao et al., 2023; Shinn et al., 2023). This is crucial as each action’s consequence, affecting the user’s information state or the environment’s state, provides vital feedback. The Planning Module uses this feedback to iteratively adjust the agent’s policy and refine future actions.
@@
  • Embodied Navigation and Manipulation: For embodied agents or robotic control systems, the Action Module grounds high-level skills or planned actions into low-level motor commands executable in physical or virtual environments (Xi et al., 2025). This serves as an intermediary, bridging the gap between cognitive planning and physical execution.
@@
Therefore the design of a comprehensive and precise action space is crucial for building robust agents, as it standardises the translation of high-level plans into executable operations. To enhance execution accuracy, action names should semantically correspond to their behavior, and natural language descriptions should be provided to clarify action usage (Xu et al., 2024a).
@@ 2.3.1. Subjective Evaluation
Subjective evaluation is particularly effective when assessing qualitative aspects of agent performance, such as overall helpfulness or user-friendliness, where quantitative metrics are difficult to define or evaluation datasets are limited. LLM-based agents are often designed to serve humans, making subjective evaluation a critical component as it reflects human criteria. This involves human evaluators directly interacting with the system or observing its performance, and then providing judgments based on predefined criteria or overall impressions. This approach often uses structured questionnaires with Likert scales to assess qualitative aspects. For instance, Franciscatto et al. (Franciscatto et al., 2025) employed a 5-point Likert scale to evaluate visibility, support, usefulness, transparency, justification, and data integration capabilities. Similarly, Shaer et al. (Shaer et al., 2024) used Likert scales to assess generated ideas for relevance, innovation, and insightfulness by both expert and novice evaluators. Furthermore, in (Schmidgall et al., 2025) the authors conducted human surveys with PhD researchers, who rated paper quality on experimental quality, report quality, and usefulness, and also simulated peer review using NeurIPS-style scores. Subjective evaluation, while offering valuable qualitative insights, shows a series of considerable challenges. Primarily, it is resource-intensive and time-consuming, necessitating the recruitment of a sufficiently large and diverse pool of human evaluators, which can incur substantial financial costs and require extensive logistical planning for study preparation and execution. Moreover, the inherent subjectivity of human judgment introduces potential biases, influenced by individual perspectives, experiences, and cultural backgrounds. These biases can disrupt the objectivity and generalizability of evaluation results. Consequently, the interpretation and integration of subjective evaluation data require careful consideration of these inherent limitations, emphasizing the need for rigorous methodological design and transparent reporting to mitigate potential biases and enhance the reliability and validity of findings.
@@ 2.3.2. Objective Evaluation
Objective evaluation uses quantitative metrics to enable computational, and comparative analysis, thereby providing concrete and measurable assessments of agent performance. The selection of appropriate metrics is crucial for evaluating specific aspects of agent performance. For Task Completion and Efficiency, evaluation focuses on the agent’s ability to achieve defined objectives with efficient resource use. This category includes metrics such as Success Rate (Chang et al., 2024), quantifying overall goal attainment; Progress Rate (Chang et al., 2024), tracking incremental advancements in multi-turn interactions; Completion Rate (Xu et al., 2024a), measuring the proportion of completed sub-tasks; Execution Efficiency (Xu et al., 2024a), assessing action efficiency relative to sub-task completion; Cost Efficiency (Xu et al., 2024a), evaluating resource consumption through token usage; Number of Steps (Xu et al., 2024b), quantifying operational effort via LLM calls; and Cost per Instance (Xu et al., 2024b), measuring monetary expenditure for API queries. In numerous applications, Accuracy and Correctness have critical importance. Metrics within this category assess the fidelity and consistency of generated information, plans, and actions. This includes Accuracy (Wang et al., 2024c), evaluating overall output correctness; Answer F1 (Liu et al., 2023), measuring query response accuracy in knowledge graph environments; Test Coverage and Bug Detection Rate (Wang et al., 2024c), assessing the agent’s ability to generate effective test cases and identify software defects. For agents designed for conversational or interactive tasks, Dialogue and Interaction Quality is assessed using metrics such as Recency, Relevance, and Importance (Hou et al., 2024). These metrics evaluate the coherence, relevance, and meaningfulness of agent dialogues, capturing the nuances of human-agent interaction. Finally, as AI agents are increasingly deployed in critical systems, Adversarial Robustness becomes a key evaluative consideration. Metrics such as Benign Utility, Utility Under Attack, and targeted Attack Success Rate (Debenedetti et al., 2024) evaluate the agent’s resilience to adversarial attacks, ensuring security and reliability in real-world deployments.
@@
To comprehensively evaluate the capabilities of agents across diverse application domains, a range of specialised benchmarks and frameworks have been developed. These include AgentBench  (Liu et al., 2023) , which features tasks simulating real-world scenarios such as web browsing and gaming, alongside code generation and execution tasks involving operating systems, databases, and knowledge graphs. TheAgentCompany (Xu et al., 2024b) provides a self-contained environment that models a small software company, including tasks like web browsing, code writing, and inter-agent communication. DevAI (Zhuge et al., 2024) presents 55 real-world AI application development tasks curated by expert annotators, while SWE-bench (Yang et al., 2024) offers 2294 software engineering problems derived from actual GitHub issues and pull requests across 12 prominent Python repositories. Text-based game environments, such as ALFWorld (Shinn et al., 2023), and Minecraft (Wang et al., 2023), are used to evaluate language agent performance in interactive, simulated settings. WebShop (Yao et al., 2022) focuses on assessing product search and retrieval capabilities, and WebArena (Zhou et al., 2023) provides a comprehensive website environment for end-to-end agent evaluation. RoCoBench (Mandi et al., 2024) evaluates multi-agent collaboration across diverse scenarios, emphasizing communication and coordination in cooperative robotics. TravelPlanner  (Xie et al., 2024) benchmarks real-world planning capabilities in language agents, and ScienceWorld  (Wang et al., 2022) evaluates reasoning and planning abilities by posing questions designed to challenge a fifth-grade student. Finally, for evaluating safety and privacy, AgentDojo  (Debenedetti et al., 2024) measures AI agent resilience to prompt injection attacks, and PrivacyLens  (Shao et al., 2024a) quantifies potential data leakage to assess privacy norm adherence in language model agents.

### v2

@@ 2.2.3. Planning
The planning module is essential for autonomous agents, enabling them to formulate future action sequences to achieve defined goals (wang2024survey). It allows agents to move in goal-directed, multi-step problem-solving approach. The planning module combines environmental feedback, current state, desired outcomes, memory, and profile information to construct a trajectory of actions that effects a transition from the agent’s current state to a target goal state. Various methodologies have emerged for incorporating LLMs into the planning process, each representing a different approach to leveraging their capabilities:
@@
  • LLM-as-Planner directly employs the inherent reasoning abilities of LLMs to generate plans from natural language instructions (li2024lasp). This paradigm emphasizes the autonomous planning capacity of LLMs, relying on their ability to interpret and translate natural language directives into actionable plans.
@@
  • LLM-as-Facilitator uses LLMs to augment existing planning algorithms, such as classic symbolic planners. In this context, LLMs may serve as translators, converting natural language problem descriptions into formal planning languages, such as the Planning Domain Definition Language (PDDL), which are then processed by external planners (li2024embodied; wang2024survey).
@@
  • Multi-Agent Planning involves the coordination of plans and actions among multiple agents to achieve a shared objective (han2024llm). This necessitates the implementation of mechanisms for inter-agent communication, negotiation, and belief revision. Collaborative planning approaches are exemplified by frameworks such as PlanGEN (parmar2025plangen), which employs specialised LLM agents for constraint checking, verification, and selection, and Master (gan2025master), which proposes a hierarchical multi-agent framework that dynamically generates collaborating agents, validates reasoning, and adjusts confidence-based scoring using Monte Carlo Tree Search (MCTS) to enhance accuracy and efficiency.
@@
One key aspect of the planning module is task decomposition, where complex tasks are broken down into smaller, more tractable sub-problems, facilitating efficient execution. For example, PlanGEN (parmar2025plangen) is specifically engineered to augment LLMs’ capacity to generate effective natural language plans through this decomposition process. Similarly, WebPilot’s Planner module (zhang2024webpilotversatileautonomousmultiagent) initiates its operation by partitioning complex web tasks into manageable subtasks, thereby constructing a flexible, high-level plan that can adapt to the inherent uncertainties of web environments. Consequently, determining the optimal sequence of these subtasks or individual actions, becomes crucial for achieving the overarching goal. In more complex scenarios, the planning module may also necessitate the allocation of available resources across various planning components (li2024lasp). This is exemplified by the TravelPlanner benchmark (xie2024travelplanner), which evaluates agents’ ability to generate travel itineraries from user queries, demanding the navigation of extensive online resources using specialised tools and adherence to user constraints. This benchmark assesses the agent’s resource allocation across diverse tasks, including city searches, flight bookings, accommodation selection, and dining arrangements. In the current literature the Planning Module employs diverse strategies in task decomposition:
@@
  • Single-Path Planning: The agent follows one trajectory of thought and action at a time, without exploring alternative possibilities. Examples include Chain of Thought prompting, where the LLM reasons step-by-step along a linear path (wang2024survey).
@@
  • Tree-Based Planning: The agent explores multiple potential thought trajectories, organizing them into a tree structure. This allows for backtracking and evaluating different options before committing to a plan. Techniques like Tree of Thoughts and methods using Monte Carlo Tree Search (MCTS) fall under this category (li2024lasp).
@@
  • Hierarchical Planning: Involves planning at different levels of abstraction, with high-level plans broken down into more detailed, low-level actions or sub-goals (xi2025rise; zhang2024webpilotversatileautonomousmultiagent).
@@
Effective integration of feedback within the planning module significantly enhances the ability of LLM-based agents to operate in complex and dynamic environments (han2024llm). This capability allows agents to move beyond static, pre-defined plans and engage in a more adaptive and iterative problem-solving process. So we can categorise planning approaches in 2 groups:
@@
  • Planning without Feedback: The agent formulates a plan upfront and executes it sequentially without adjusting based on intermediate outcomes (han2024llm). This approach is appropriate for tasks of lower complexity, including conversational interaction.
@@
  • Planning with Feedback: The agent receives information from the environment, humans, or other models about the progress and outcomes of its actions (li2024lasp). This feedback is then used to dynamically adjust and refine the plan. Environmental Feedback can include task completion signals or observations after taking an action (yao2023react). Human feedback can provide guidance or corrections (franciscatto2025cbr). Model-based feedback, such as self-reflection or critique from another AI model, can also be valuable (shinn2023reflexion).
@@
Finally, careful consideration of memory and profiling is vital for effective planning. The memory module provides the historical context, past experiences, and relevant knowledge needed to formulate informed plans. For instance, past successes or failures in similar situations, user preferences, or environmental observations stored in memory can guide the planning process (shinn2023reflexion). The agent’s profile, defined by the profiling module, can also influence the style and focus of the planning process. For example, an agent profiled as risk-averse might prioritise plans with higher certainty of success, while a creative problem-solver profile might encourage the exploration of novel or less conventional plans (zhang2024agentic).
@@ 2.2.4. Action
The primary goal of the Action Module is to translate the agent’s formulated plans into specific outcomes within the environment. It acts as the execution engine, directly engaging with the surrounding world (wang2024survey). The intended outcomes of agent actions are diverse and intrinsically linked to the agent’s designated task. Actions may be directed towards achieving concrete objectives, such as item crafting in a game (wang2023voyager) or software development task completion (wang2024openhands), where each action contributes directly to the final goal. In the context of a web browsing agent (lai2024autowebglm), actions can include navigating or interacting to specific pages. Actions may also facilitate information sharing and collaboration with other agents (gan2025master) or human users  (franciscatto2025cbr), including conversational exchanges and feedback provision. Furthermore, agents perform exploratory actions to gain new knowledge and expand perception (wen2024autodroidv1), optimizing learning and performance through exploration-exploitation. The Action Module receives from the Planning Module a sequence of steps or a specific action to be taken and translates them into executable actions (wang2024survey). The Action Module’s operational mechanism varies with the agent’s architecture and action space, which collectively define the agent’s interaction capabilities. Actions can be broadly categorised as follows:
@@
  • External Tool and API Interaction: LLM-driven agents frequently generate natural language commands or structured calls to external tools and APIs (schick2023toolformer). This allows them to interact with the external world, using resources like search engines for web navigation (lai2024autowebglm) (including clicking, form filling, and scrolling), databases via SQL queries (shen2024demonstration), code execution environments (wang2024openhands), PDDL solver execution (chang2024agentboard), or specific application such as FlightSearch API with correctly formatted parameters (xie2024travelplanner). To simplify the integration of diverse APIs, the Model Context Protocol (MCP) (unknown-author-2024) standardises API access, eliminating manual tracking and description. MCP’s Host-Client-Server architecture facilitates dynamic tool discovery and execution, enabling any MCP-compatible LLM application to use connected server tools.
@@
  • Internal Response Generation: This category includes actions that generate internal responses, such as providing explanations or refining plans (yao2023react; shinn2023reflexion). This is crucial as each action’s consequence, affecting the user’s information state or the environment’s state, provides vital feedback. The Planning Module uses this feedback to iteratively adjust the agent’s policy and refine future actions.
@@
  • Embodied Navigation and Manipulation: For embodied agents or robotic control systems, the Action Module grounds high-level skills or planned actions into low-level motor commands executable in physical or virtual environments (xi2025rise). This serves as an intermediary, bridging the gap between cognitive planning and physical execution.
@@
Therefore the design of a comprehensive and precise action space is crucial for building robust agents, as it standardises the translation of high-level plans into executable operations. To enhance execution accuracy, action names should semantically correspond to their behavior, and natural language descriptions should be provided to clarify action usage (xu2024crab).
2.2.5. Agent Topologies
The design strategies employed for LLM Agents vary significantly based on the complexity and domain of the target tasks. They often involve augmenting the LLM core with specialized architectures for memory management, external knowledge access, and collaborative reasoning:
  • Single Agent + action tools: This strategy equips a single LLM agent with the ability to execute code or commands within a dedicated and isolated environment, typically referred to as a sandbox or a Graphical User Interface (GUI), which provide a set of tools for the interaction with operating systems or code generation (zhuge2024agent; shinn2023reflexion; lai2024autowebglm). Usually a sandbox is provided for a secure execution space for arbitrary code generated by the agent, mitigating potential safety risks associated with deploying unpredictable LLM output directly into production systems.
  • Agent + Information Retrieval (RAG, Tools, APIs): This approach focuses on overcoming the limitations of an LLM’s static internal knowledge (hallucination, outdated facts) by augmenting its input context with external, non-parametric knowledge retrieved from diverse sources (li2025search; wu2024avatar; zhang2024agentic). This approach focuses on overcoming the limitations of an LLM’s static internal knowledge (hallucination, outdated facts) by augmenting its input context with external, non-parametric knowledge retrieved from diverse sources. This technique is used for factual and updated response such as in medical and enterprise domains.
  • Multi-Agent Systems (MAS): This paradigm overcomes the cognitive bottlenecks of monolithic LLMs by distributing complex problem-solving across a network of specialized agents. Rather than relying on a single inference chain, MAS architectures enforce coordination and communication to achieve shared objectives. Interaction patterns typically fall into two categories: collaborative execution, where agents assume distinct functional roles (e.g., Planner, Executor, Reviewer) to parallelize sub-tasks (erdogan2025plan; gebreab2024llm); and iterative refinement, where agents engage in debate or feedback loops to critique and improve outputs before final delivery (chan2023chateval; chen2025improving). This topology is increasingly essential in high-stakes domains, such as clinical trial analysis (yue2024clinicalagent) and autonomous software maintenance (tao2024magis).
Table 1. Key LLM agent frameworks and benchmarks. The benchmarks are grouped by their primary domain focus to highlight differences in environments and key characteristics.
Benchmark/Framework Evaluation Focus Environment
AgentBench (liu2023agentbench) Multi-domain capabilities OS, databases, knowledge graphs, web browsers.
TheAgentCompany (xu2024theagentcompany) Simulated professional software development work Self-contained simulated corporate environment (OS, web tools, worker communications).
DevAI (zhuge2024agent) AI application development tasks Coding environment.
SWE-bench (yang2024swe) Real-world software engineering maintenance 12 Python GitHub repositories.
WebShop (yao2022webshop) E-commerce product search and retrieval Simulated online shopping website.
WebArena (zhou2023webarena) End-to-end general web task completion Comprehensive, fully functional website sandboxes.
ALFWorld (shridhar2020alfworld) Interactive situated decision-making Text-based game environment.
Minecraft (wang2023voyager) Open-ended exploration and complex task execution 3D voxel-based game world.
RoCoBench (mandi2024roco) Multi-agent collaboration and coordination Cooperative robotics scenarios.
TravelPlanner (xie2024travelplanner) Real-world constraint satisfaction planning Information travel scenarios.
ScienceWorld (wang2022scienceworld) Scientific reasoning and procedural planning Simulated science laboratory environment.
AgentDojo (debenedetti2024agentdojo) Safety and adversarial resilience Adversarial prompt injection setups.
PrivacyLens (shao2024privacylens) Privacy norm adherence and data protection Scenarios involving sensitive data handling.
@@ 2.3.1. Subjective Evaluation
Subjective evaluation is particularly effective when assessing qualitative aspects of agent performance, such as overall helpfulness or user-friendliness, where quantitative metrics are difficult to define or evaluation datasets are limited. LLM-based agents are often designed to serve humans, making subjective evaluation a critical component as it reflects human criteria. This involves human evaluators directly interacting with the system or observing its performance, and then providing judgments based on predefined criteria or overall impressions. This approach often uses structured questionnaires with Likert scales to assess qualitative aspects. For instance, Franciscatto et al. (franciscatto2025cbr) employed a 5-point Likert scale to evaluate visibility, support, usefulness, transparency, justification, and data integration capabilities. Similarly, Shaer et al. (shaer2024ai) used Likert scales to assess generated ideas for relevance, innovation, and insightfulness by both expert and novice evaluators. Furthermore, in (schmidgall2025agent) the authors conducted human surveys with PhD researchers, who rated paper quality on experimental quality, report quality, and usefulness, and also simulated peer review using NeurIPS-style scores. Subjective evaluation, while offering valuable qualitative insights, shows a series of considerable challenges. Primarily, it is resource-intensive and time-consuming, necessitating the recruitment of a sufficiently large and diverse pool of human evaluators, which can incur substantial financial costs and require extensive logistical planning for study preparation and execution. Moreover, the inherent subjectivity of human judgment introduces potential bias, influenced by individual perspectives, experiences, and cultural backgrounds. These biases can disrupt the objectivity and generalizability of evaluation results. Consequently, the interpretation and integration of subjective evaluation data require careful consideration of these inherent limitations, emphasizing the need for rigorous methodological design and transparent reporting to mitigate potential biases and enhance the reliability and validity of findings.
@@ 2.3.2. Objective Evaluation
Objective evaluation uses quantitative metrics to enable computational, and comparative analysis, thereby providing concrete and measurable assessments of agent performance. The selection of appropriate metrics is crucial for evaluating specific aspects of agent performance. For Task Completion and Efficiency, evaluation focuses on the agent’s ability to achieve defined objectives with efficient resource use. This category includes metrics such as Success Rate (chang2024agentboard), quantifying overall goal attainment; Progress Rate (chang2024agentboard), tracking incremental advancements in multi-turn interactions; Completion Rate (xu2024crab), measuring the proportion of completed sub-tasks; Execution Efficiency (xu2024crab), assessing action efficiency relative to sub-task completion; Cost Efficiency (xu2024crab), evaluating resource consumption through token usage; Number of Steps (xu2024theagentcompany), quantifying operational effort via LLM calls; and Cost per Instance (xu2024theagentcompany), measuring monetary expenditure for API queries. In numerous applications, Accuracy and Correctness have critical importance. Metrics within this category assess the fidelity and consistency of generated information, plans, and actions. This includes Accuracy (wang2024survey), evaluating overall output correctness; Answer F1 (liu2023agentbench), measuring query response accuracy in knowledge graph environments; Test Coverage and Bug Detection Rate (wang2024survey), assessing the agent’s ability to generate effective test cases and identify software defects. For agents designed for conversational or interactive tasks, Dialogue and Interaction Quality is assessed using metrics such as Recency, Relevance, and Importance (hou2024my). These metrics evaluate the coherence, relevance, and meaningfulness of agent dialogues, capturing the nuances of human-agent interaction. Finally, as AI agents are increasingly deployed in critical systems, Adversarial Robustness becomes a key evaluative consideration. Metrics such as Benign Utility, Utility Under Attack, and targeted Attack Success Rate (debenedetti2024agentdojo) evaluate the agent’s resilience to adversarial attacks, ensuring security and reliability in real-world deployments.
@@
To comprehensively evaluate the capabilities of agents across diverse application domains, a range of specialised benchmarks and frameworks have been developed. These include AgentBench  (liu2023agentbench) , which features tasks simulating real-world scenarios such as web browsing and gaming, alongside code generation and execution tasks involving operating systems, databases, and knowledge graphs. TheAgentCompany (xu2024theagentcompany) provides a self-contained environment that models a small software company, including tasks like web browsing, code writing, and inter-agent communication. DevAI (zhuge2024agent) presents 55 real-world AI application development tasks curated by expert annotators, while SWE-bench (yang2024swe) offers 2294 software engineering problems derived from actual GitHub issues and pull requests across 12 prominent Python repositories. Text-based game environments, such as ALFWorld (shridhar2020alfworld), and Minecraft (wang2023voyager), are used to evaluate language agent performance in interactive, simulated settings. WebShop (yao2022webshop) focuses on assessing product search and retrieval capabilities, and WebArena (zhou2023webarena) provides a comprehensive website environment for end-to-end agent evaluation. RoCoBench (mandi2024roco) evaluates multi-agent collaboration across diverse scenarios, emphasizing communication and coordination in cooperative robotics. TravelPlanner  (xie2024travelplanner) benchmarks real-world planning capabilities in language agents, and ScienceWorld  (wang2022scienceworld) evaluates reasoning and planning abilities by posing questions designed to challenge a fifth-grade student. Finally, for evaluating safety and privacy, AgentDojo  (debenedetti2024agentdojo) measures AI agent resilience to prompt injection attacks, and PrivacyLens  (shao2024privacylens) quantifies potential data leakage to assess privacy norm adherence in language model agents. All these framework and benchmarks are summarised in Table 1.